OpenStudy (anonymous):

solve the system using matrices (Gaussian elimination method). if there is no solution or if the systems equation dependent, so state ? x + y + z = 9 x - y + 5z = 17 2x + y + z = 13

OpenStudy (amistre64):

matrix it up and row reduce ....

OpenStudy (amistre64):

if you get a row that is inconsistent; its bad; otherwise there are solutions

OpenStudy (anonymous):

how

OpenStudy (amistre64):

how what? how to you form a matrix? how do you perform elementary matrix operations? how do you ... what?

OpenStudy (amistre64):

i assume since you ask the question that your familiar with at least some of these processes and are not actually wanting to know all the stuff about matrices

OpenStudy (anonymous):

well this is a take home test, so i dont wana give it to her and its wrong answer

OpenStudy (anonymous):

im tring to get the answer so i compare it to mine wen i try attempt to solve it

OpenStudy (amistre64):

there are 3 elementary row operations that can be done on a matrix; do you recall what they are?

OpenStudy (anonymous):

yeahh

OpenStudy (amistre64):

if we know those operations we can then strip all the stuff from our equations and keep only numbers and signs: x + y + z = 9 x - y + 5z = 17 2x + y + z = 13 becomes the matrix 1 1 1 9 1 -1 5 17 2 1 1 13

OpenStudy (amistre64):

what are our 3 elementary row operations that we can do?

OpenStudy (anonymous):

yep i did that arledyy

OpenStudy (anonymous):

, -1 5 1 1

OpenStudy (anonymous):

-1(5) + 1(1)

OpenStudy (amistre64):

you trying to cramer this thing up? thats madness

OpenStudy (anonymous):

oh my bad lol, sorry got confused lol

OpenStudy (amistre64):

1 1 1 9 1 -1 5 17 ; *-1 2 1 1 13 ; /-2 1 1 1 9 -1 1 -5 -17 -1 -1/2 -1/2 -13/2 1 1 1 9 0 2 -4 -8 0 1/2 1/2 5/2 1 1 1 9 0 2 -4 -8 ; /2 0 1/2 1/2 5/2 ; *-2 1 1 1 9 0 1 -2 -4 0 -1 -1 -5 1 1 1 9 0 1 -2 -4 0 0 -3 -9 ; /-3 1 1 1 9 0 1 -2 -4 0 0 1 3 well, the good noews is its consistent

OpenStudy (anonymous):

wow thanks man, thats the hole thing ?

OpenStudy (amistre64):

you really dont know whats going on with this do you ....

OpenStudy (amistre64):

thats just half of it

OpenStudy (anonymous):

honestly i have no idea man. im not gona lie and say i do.

OpenStudy (anonymous):

this is a take home test and im tring to get this right.

OpenStudy (amistre64):

the matrix method is not a good way to do by hand; elimination and substitution are the way to go if your doing it by hand

OpenStudy (amistre64):

so far, all ive done is get this to a row eschelon form; I still have to reduce it by getting 0s above all my 1s to read off the unique solution on the right side

OpenStudy (anonymous):

im so lost lol. im tring to rite this down and im like 0.0

OpenStudy (amistre64):

1 1 1 9 0 1 -2 -4 0 0 1 3 .... z = 3 1 1 1 9 0 1 0 -4 -2(3) = -10; y=-10 0 0 1 3 .... z = 3 1 0 0 9 -(3)+(10) = 16 0 1 0 -10 ... y=-10 0 0 1 3 .... z = 3 1 0 0 16 ... x = 16 0 1 0 -10 ... y=-10 0 0 1 3 .... z = 3 now the best thing to do is dbl chk it with something that can handle the repetitive nature; like a computer :)

OpenStudy (amistre64):

rref[1, 1, 1, 9;1, -1, 5, 17;2, 1, 1, 13] lets see if the wolf like this :)

OpenStudy (anonymous):

lmao, see if u did mma, or hit the gymm with me i woulda helped u lmao. but as u can see im retricein this.

OpenStudy (anonymous):

i ment im garabge lol

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=rref%7B%7B1%2C+1%2C+++1%2C+++9%7D%2C%7B1%2C++-1%2C++5%2C++17%7D%2C%7B2%2C+++1%2C+++1%2C++13%7D%7D wolf like a different format; but it says I missed the mathing someplace

OpenStudy (amistre64):

1 1 1 9 0 1 -2 -4 0 0 1 3 ; *2, *-1 1 1 1 9 0 0 -1 -3 ----------------- 1 1 0 6 0 1 -2 -4 0 0 2 6 ----------------- 0 1 0 2 1 1 0 6 0 1 0 2 :*-1 0 0 1 3 1 1 0 6 0 -1 0 -2 -------------- 1 0 0 4 1 0 0 4 0 1 0 2 0 0 1 3 so our solutions are: x=4, y=2, amd z=3

OpenStudy (amistre64):

thatll teach me to try something new lol

OpenStudy (anonymous):

lmao i dont even noe wat to rite to show work. LOL

OpenStudy (amistre64):

and unless you know what im doing, this will make absolutely no sense ... :)

OpenStudy (amistre64):

look thru your material to see how it shows the row operations and see if you can line some things up according to the format your material provides you with

OpenStudy (anonymous):

okk ima chek but can u stick around to help me ? you will be my life savor lol.

OpenStudy (amistre64):

i got school work to catch up on too, but if im around ill see what i can do within reason

OpenStudy (anonymous):

okaay thankss lovee. big help !

OpenStudy (amistre64):

as i said prior; this question itself presupposes that you have some idea as to whats going on. otherwise there is too much behind the scenes work to try to explain

OpenStudy (anonymous):

yeah i noe im looking at it now, but thanks alot. im cheking and reviewing it now